OpenStudy (anonymous):

a regular polygon A has the midpoints of its edges joined to form a smaller hexagon B to get a third hexagon C. What is the ratio of the area of hexagon C to the area of hexagon A?

OpenStudy (jack1):

equation for area of a hexagom =...?

OpenStudy (jack1):

Area of Polygon = ¼ × n × (Side length)^2 / tan(π/n) n = number of sides = 6 for a hexagon so: \[Area.of.Hexagon = \frac 14 × 6 × (side)^2 / \tan(π/6)\]

OpenStudy (jack1):

\[Area.of.Hexagon = 1.5 × (side)^2 / \tan(π/6)\] now tan (pi/6) = 1/sqrt 3 so \[Area.of.Hexagon = 1.5 × \sqrt 3 × (side)^2 \]

OpenStudy (jack1):



OpenStudy (jack1):

 ratio of side length = 2:1 new sidelength is 1/2 original side length so new area is (1/2)^2=1/4 so do that again new side length = 1/4 of original so new area is (1/4)^2=1/16 th area of original area

Directrix (directrix):

Similar problem: http://openstudy.com/study#/updates/507d2c0ae4b040c161a2bc42

OpenStudy (jmark):

Suppose The Length Of One Side Of The Hexagon A Is 1cm (Interior Angle Of A Regular Hexagon Is 120 Degrees). Midpoint Would Be 0.5 cm Away From The Edge. Then for Hexagon B, The Length Of One Side Using The Cosine Rule. =>Cos120 = (0.5^2 + 0.5^2 -c^2)/2*0.5*0.5 =>c = 0.866cm The Midpoint For ""B"" Would Be 0.866/2 cm From The Edge Of ""B"" which is 0.433cm Same For Hexagon C =>Cos120 = (0.433^2 + 0.433^2 -x^2)/2*0.433*0.433 =>x = 0.75 cm The Hexagons (A, B & C) Are Geometrically Similar we have =>(length1/length2)^2 = Area1/Area2 hence =>(0.75/1)^2 = Area C/Area A we have 1cm was the length of one side of A and 0.75 was the length of one side of C =>Area C/ Area A = 0.5625 = 9/16 ratio of the area of hexagon C to the area of hexagon A =9/16 Know more about Perimeter of a Polygon http://goo.gl/HN2M6u